What number of liters of STP CO2 is produced when 7x10^23 molecules of O2 is consumed? Show the solution

1
Expert's answer
2022-05-31T20:23:04-0400

n(O2) = N/Na

n(O2) = 7×1023/6.022×1023 = 1.162 mole

V(O2) = n•Vm

V(O2) = 1.162•22.4 = 26 litres